### v1

# mO shizhe[#1]

Abstract
Retrieval-augmented generation (RAG) has become a main technique for alleviating hallucinations in large language models (LLMs). Despite the integration of RAG, LLMs may still present unsupported or contradictory claims to the retrieved contents. In order to develop effective hallucination prevention strategies under RAG, it is important to create benchmark datasets that can measure the extent of hallucination. This paper presents RAGTruth, a corpus tailored for analyzing word-level hallucinations in various domains and tasks within the standard RAG frameworks for LLM applications. RAGTruth comprises nearly 18,000 naturally generated responses from diverse LLMs using RAG. These responses have undergone meticulous manual annotations at both the individual cases and word levels, incorporating evaluations of hallucination intensity. We not only benchmark hallucination frequencies across different LLMs, but also critically assess the effectiveness of several existing hallucination detection methodologies. Furthermore, we show that using a high-quality dataset such as RAGTruth, it is possible to finetune a relatively small LLM and achieve a competitive level of performance in hallucination detection when compared to the existing prompt-based approaches using state-of-the-art large language models such as GPT-4.111The RAGTruth dataset is available at https://github.com/ParticleMedia/RAGTruth
1 Introduction
Large language models (LLMs) have achieved remarkable success in a variety of tasks, including text generation Li et al. (2022), machine translation Kocmi and Federmann (2023), and question answering Zhao et al. (2023). However, one of the key challenges in deploying LLMs in real-world applications is their tendency to hallucinate Kaddour et al. (2023). Hallucination in the context of LLMs usually refers to a situation where the model generates content that is not based on factual or accurate information Rawte et al. (2023). The occasional generation of outputs that appear plausible but are factually incorrect significantly undermine the reliability of LLMs in real-world scenarios, such as medical diagnoses Pal et al. (2023) and news summarization Shen et al. (2023).
To reduce hallucination, various methods have been developed that can be applied at different stages of LLM lifecycle, including pre-training Brown et al. (2020), supervised fine-tuning Zhou et al. (2023); Zhang et al. (2023a), RLHF Ouyang et al. (2022); Lin et al. (2022), and inference Dhuliawala et al. (2023); Gao et al. (2023). In terms of detection, methods are developed by examining the model’s intrinsic state Guo et al. (2017), comparing it with external data and tools Chern et al. (2023), or leveraging the LLM’s inherent powerful capabilities for self-checking Agrawal et al. (2023); Manakul et al. (2023). RAG is extensively used to supply LLMs with updated, relevant knowledge, significantly mitigating hallucination. Nevertheless, even with RAG and other enhancements, LLMs still produce statements that are either unfounded or contradict the information provided in the retrieved references Shuster et al. (2021).
Despite the growing awareness of the hallucination phenomenon, the understanding of hallucination in LLMs is still in its early stages. One key challenge is the lack of high-quality, large-scale datasets specifically designed for hallucination detection. This issue is particularly acute in RAG settings. Due to the relatively low hallucination ratio, a substantial increase in annotation resources is needed. Existing datasets for LLM hallucination detection are predominantly synthesized Li et al. (2023). For instance, in Liu and Liu (2023); Longpre et al. (2021) prompts conflicting with conventional knowledge are purposely generated to trigger hallucinations. While these approaches are efficient at generating hallucinations, the resulting artificial hallucinations can substantially differ from those that naturally occur. In Chen et al. (2023); Hu et al. (2023), hallucination datasets are developed by manual annotations of naturally produced LLM responses. However, these datasets are of limited size and are not specifically focused on the RAG scenario.
In this paper, we introduce a large-scale high-quality dataset specifically designed for word-level hallucination detection for RAG applications. Using this dataset, we have conducted an extensive benchmarking of mainstream LLMs to assess their tendency to generate hallucinations, as well as evaluate current methods for hallucination detection. Additionally, we have demonstrated superior performance in identifying hallucinations by fine-tuning LLM with RAGTruth dataset. Our key contributions are:
  1. (i)
    We propose RAGTruth, a large-scale word-level hallucination evaluation dataset specifically for the RAG scenario across several common tasks. It consists of nearly 18,000 fully annotated natural responses generated from major open-source and closed-source LLMs.
  2. (ii)
    We perform a comprehensive comparison of different hallucination detection methods at both the passage and word levels.
  3. (iii)
    We present a simple baseline method of fine-tuning LLM for hallucination detection. It is shown that by fine-tuning the Llama-2-13B model on the RAGTruth training data, we can achieve results competitive to the existing prompt-based approaches using GPT-4. This shows the potential of developing better hallucination detection methods using RAGTruth.
  4. (iv)
    We show that by using our finetuned hallucination detector, it is possible to significantly reduce the occurrence of hallucinations in the responses from LLMs. The improvement holds even for models with inherently low hallucination rates, such as GPT-4.
2 Related Work
2.1 Hallucination of Large Language Models
In Zhang et al. (2023b), the hallucination of large language models (LLMs) is characterized into three categories: 1) input-conflicting hallucination 2) context-conflicting hallucination 3) fact-conflicting hallucination. Existing research has demonstrated that incorporating up-to-date, relevant knowledge in the prompt can effectively reduce fact-conflicting hallucination Vu et al. (2023); Lewis et al. (2021). This approach, referred to as Retrieval-Augmented Generation (RAG), is widely used in real-world LLM applications. For instance, Google Bard  [^2] https://bard.google.com and Microsoft BingChat  [^3] https://www.bing.com have implemented this technique. Consequently, the detection and mitigation of input-conflicting and context-conflicting hallucinations within the RAG framework are becoming increasingly important.
2.2 Hallucination Detection Methods
Researchers have been exploring various methods to enhance the reliability of LLMs by detecting hallucinations. In Azaria and Mitchell (2023); Xiao and Wang (2021); Malinin and Gales (2021), intrinsic model uncertainty metrics such as token-level probability and entropy are used to detect hallucinations. When direct access to output uncertainty is not feasible, as in the case with limited APIs like GPT-4, an alternative approach involves employing a fully accessible LLM as a proxy Manakul et al. (2023). In Falke et al. (2019); Barrantes et al. (2020), natural language inference modules are adapted to check the information consistency between the articles and their summaries, and it has been shown that external knowledge is helpful for detecting factual hallucinations. Guo et al. (2022); Mallen et al. (2022). Additionally, methods that harness the inherent capabilities of LLMs have been proposed, such as verbalize-based and consistency-based methods Xiong et al. (2023); Manakul et al. (2023). These techniques aim to detect hallucinations without relying on internal states or external data and tools.
Figure 1: Data gathering pipeline. Taking a data-to-text task as an example, our data gathering pipeline includes 2 steps: 1) response generation. We generated responses with multiple LLMs and natural prompts. 2) human annotation. Human labeler annotated hallucinated spans in LLM responses.
2.3 Hallucination Evaluation Datasets
Constructing a hallucination dataset can be a labor-intensive task, as it necessitates the manual identification of ungrounded or falsely claimed statements in the output of LLMs. Such a dataset can be particularly valuable in the context of RAG scenarios, owing to its significantly lower incidence of hallucinations.
Hallucinate datasets largely fall into two categories: synthetic dataset and manually-crafted dataset. Synthetic datasets typically involve purposely triggering Language Model to produce spurious responses Li et al. (2023) or simply inserting knowledge conflicts into the output Longpre et al. (2021). While these methods are effective in generating a substantial volume of hallucinated responses, these responses may not follow the same distribution as the naturally generated responses. On the other hand, manually annotated datasets offer a closer representation of the LLMs’ true generative capabilities Vu et al. (2023); Lewis et al. (2021). However, the considerable labor and expense required for annotation lead to datasets of limited size and length and inconsistent quality. Specifically, there is a lack of manually curated datasets of substantial size and quality to facilitate hallucination detection in RAG scenarios.
3 Construction Process of RAGTruth
We established a data generation and annotation pipeline as shown in Figure 1.
3.1 Response Generation
Tasks and Data Sources
We selected three widely recognized RAG tasks for response generation: Summarization, Question Answering, and Data-to-Text generation.
For the summarization task, we randomly selected documents from the training set of the well-known CNN/Daily Mail dataset See et al. (2017) as well as recent news articles from a prestigious news platform. LLMs were prompted to generate summaries for each of the source documents.
For the task of question answering, we conducted a random sampling from the training set of MS MACRO Nguyen et al. (2016). To reduce the annotation cost, only 3 retrieved passages for each question were preserved. Then we prompted LLMs to generate answers for each question solely based on the retrieved passages.
For the data-to-text generation task, we randomly sampled businesses in the restaurant and nightlife categories from the Yelp Open Dataset Yelp (2021). This dataset served as our structured data source. To streamline the annotation process, we focused only on the following business information fields: BusinessParking, RestaurantsReservations, OutdoorSeating, WiFi, RestaurantsTakeOut, RestaurantsGoodForGroups, Music, and Ambience. We then prompted LLMs to generate an objective text introduction for each business.
Reference: Passage 1:1 If you are just interested in checking the gender, you can go as early as 17 weeks. 2 Any earlier, and the results may not be accurate, leading to an extra coat of paint for your bedroom. 3 To get good pictures, 3D ultrasounds are best performed between 20 and 32 weeks, and the best pictures are between 24 and 30 weeks.
passage 2:…
passage 3:…
Question: how to prepare to get an ultrasound?
Response: Based on the given passages, here’s how to prepare for an ultrasound:* For a gender reveal ultrasound, wait until at least 17 weeks of pregnancy to ensure accurate results. * For 3D ultrasounds, schedule the appointment between 20 and 32 weeks of pregnancy for the best pictures
Hallucination or Not: Yes
Span: between 20 and 32 weeks of pregnancy for the best pictures
Type: Evident Conflict
Table 1: An example of our annotated data from Question Answering task. It contains reference, question, model generated response with hallucination label, annotated hallucination span and corresponding reason.
Models
The following six models with strong instruction-following ability are used for response generation: GPT-3.5-turbo-0613 and GPT-4-0613 from OpenAI OpenAI (2023); Mistral-7b-Instruct from Mistral AI Jiang et al. (2023); Llama-2-7B-chat, Llama-2-13B-chat and Llama-2-70B-chat (4bit quantized) [^4] https://huggingface.co/TheBloke/Llama-2-70B-Chat-AWQ from Meta Touvron et al. (2023). To ensure a fair comparison, the prompts used for response generation are kept straightforward with subtle differences among various models to optimize their performance. We provide detailed prompts in the Appendix B.
For each sample, we collected one response from each model. As a result, we got a total of 6 responses for each input sample.
3.2 Human Annotation
Identifying AI-generated hallucinations is a challenging task. It requires a strong capacity for critical thinking to understand the logical flow of various texts, along with meticulous attention to detail for spotting subtle inaccuracies and inconsistencies. Moreover, a certain level of media literacy and knowledge of current affairs is crucial to grasp the subjects discussed in news-related sample data. Therefore, we chose annotators who are proficient in English and possess a bachelor’s degree in relevant fields to ensure the accuracy and reliability of the annotation results.
The annotators are invited to perform annotation tasks using Label Studio Tkachenko et al. (2020-2022). Each labeling task is presented within one page, comprising the following components: 1) The original input text provided to the AI models; 2) A set of 6 responses, generated by different AI models.
Their task was to annotate the specific spans of the generated text that contains hallucinated information and categorize them into four types, based on the criteria outlined below. For examples of each type, please refer to Appendix A.
Evident Conflict:
for when generative content presents direct contraction or opposition to the original input. These conflicts are easily verifiable without extensive context, often involving clear factual errors, misspelled names, incorrect numbers, etc.
Subtle Conflict:
for when generative content presents a departure or divergence from the original input, altering the intended contextual meaning. These conflicts often involve substitution of terms that carry different implications or severity, requiring deeper understanding of their contextual applications.
Evident Introduction of Baseless Information:
for when generated content includes information not substantiated in the original input. It involves the creation of hypothetical, fabricated, or hallucinatory details lacking evidence or support.
Subtle Introduction of Baseless Information:
is when generated content extends beyond the original input by incorporating inferred details, insights, or sentiments. This additional information lacks verifiability and might include subjective assumptions or commonly observed norms rather than explicit facts.
3.3 Annotations for Adaptive Evaluation
Hallucination annotation is not always straightforward. In contentious cases, additional annotations are provided to accurately reflect these situations. This approach enables users to adopt various evaluation strategies tailored to their specific application circumstances.
Incorrectly Refusing to Answer
In the question-answering task, the model is required to decline to answer when the context lacks relevant information. However, there are instances where the model mistakenly refuses to answer despite the presence of the relevant context. These refusal responses are not labeled as hallucinations but are instead marked as incorrect_refusal.
Task # Instance # Resp. Context Length Resp. Length Hallucination
Mean Max Mean Max # Resp. % Resp. # Span
Summarization(CNN/DM) 628 3768 648 1749 124 632 1230 32.6% 1570
Summarization(Recent News) 316 1896 369 481 89 240 540 28.5% 613
Question Answering 992 5952 243 509 119 381 1844 31.0% 3155
Data-to-text 1037 6222 354 1253 159 369 ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
Overall 2973 17838 380 1749 131 632 7918 44.4% 14781
Table 2: The basic statistics of RAGTruth. $\dagger$: The data presented in the table above include hallucination spans caused by null values in JSON. If these spans are removed, the hallucination response number, hallucination response rate, and hallucination span number of date-to-text tasks are 3898, 62.6%, and 7797, respectively.
Differences in Handling Null Value
In the data-to-text task, certain fields sometimes are with null values. We observed that in the generated results, null is often interpreted as false by some models. Since the more common expressions for negation in our dataset are the boolean value False or the text No, we labeled these instances as hallucinations (evident introduction of baseless info), and provided a special span-level annotation named due_to_null. Users can independently choose whether to consider these spans as hallucinations. Additionally, we have included an option in the evaluation script that allows users to decide whether to include these spans in the evaluation.
Figure 2: Frequency of different types of hallucination by task.
4 Hallucination Benchmark Analysis
4.1 Basic Statistics
We presented detailed statistics of RAGTruth in Table 2. Compared to existing datasets for hallucination detection Cao et al. (2023); Kamoi et al. (2023), the RAGTruth dataset is considerably large in scale. The corpus contains a total of 2,973 instances of data, which include 944 instances for summarization, 992 instances for question answering, and 1,037 instances for data-to-text generation. Each instance comprises responses from 6 different models. As shown in Table 2, the RAGTruth dataset also features longer prompt and response lengths than existing datasets for hallucination detection Wang et al. (2020).
4.2 Hallucination Statistics
Hallucination Types
As shown in Figure 2, the generation of information baseless in the context was significantly more prevalent than the generation of information conflicting with the context. Within the two major categories of baseless info and conflict, the more severe hallucinations, namely Evident baseless info and Evident conflict, respectively, account for a significant portion. This observation highlights the importance and challenges of hallucination mitigation, even in RAG settings.
Hallucination vs Tasks
As shown in Table 2, across the three tasks, the data-to-text task exhibited the highest frequency of hallucinations in its responses. When processing Yelp business data, the models often generated information not present in the input, influenced by stereotypes inherent in the business data. Furthermore, inconsistent handling of JSON format data also contributed to a significant number of hallucinations. In contrast, the hallucination rate for the summarization task was relatively lower. Interestingly, the models did not show a higher rate of hallucinations for recent news compared to outdated news. This could be attributed to the shorter context length in the recent news subtask compared to the CNN/DM subtask.
Model Summarization Question Answering Data-to-text Overall
# Resp. # Span Density # Resp. # Span Density # Resp. # Span Density # Resp. # Span
GPT-3.5-turbo-0613 64 74 0.07 81 97 0.13 283 397 0.18 428 568
GPT-4-0613 78 85 0.08 48 57 0.06 293 359 0.27 419 501
Llama-2-7B-chat 451 542 0.61 542 1082 0.62 899 1812 1.30 1892 3436
Llama-2-13B-chat 333 395 0.47 426 706 0.51 988 2840 1.54 1747 3941
${}^{\dagger}$ 220 254 0.27 354 583 0.44 873 1861 1.17 1447 2698
Mistral-7B-Instruct 624 842 0.86 393 630 0.61 968 2174 1.53 1985 3646
Table 3: Hallucination counts and density of models. $\dagger$: We used 4-bit quantized version of Llama-2-70B-chat.
CLB Summarization Data2text QA
1 $0.31_{(102,317]}$ $1.55_{(117,237]}$ $0.55_{(79,163]}$
2 $0.33_{(317,387]}$ $1.49_{(237,290]}$ $0.45_{(163,196]}$
3 $0.38_{(387,481]}$ $1.54_{(290,355]}$ $0.53_{(196,266]}$
4 $0.41_{(481,810]}$ $1.49_{(355,448]}$ $0.58_{(266,327]}$
5 $0.50_{(810,1749]}$ $1.53_{(448,1253]}$ $0.54_{(327,509]}$
RLB Summarization Data2text QA
1 $0.36_{(7,74]}$ $1.24_{(35,120]}$ $0.14_{(1,70]}$
2 $0.34_{(74,92]}$ $1.21_{(120,136]}$ $0.26_{(70,98]}$
3 $0.33_{(92,110]}$ $1.63_{(136,161]}$ $0.38_{(98,127]}$
4 $0.48_{(110,144]}$ $1.85_{(161,204]}$ $0.67_{(127,166]}$
5 $0.42_{(144,632]}$ $1.67_{(204,369]}$ $1.21_{(166,381]}$
Table 4: Average number of hallucinations per response in different context length buckets (CLB) and response length buckets (RLB) for the three types of tasks. The subscript denotes the minimum and maximum length of this bucket.
Hallucination vs Models
Table 3 illustrates that among the data we collected, OpenAI’s two models demonstrated notably lower hallucination rates compared to others. Specifically, GPT-4-0613 exhibited the lowest hallucination frequency.
To more clearly compare the hallucination rate of different models, we calculated the hallucination density for each model across three tasks. Hallucination density is defined as the average number of hallucination spans per hundred words in a response. In the Llama2 series, a clear negative correlation was observed between the model scale and the hallucination density. Despite its strong performance in various benchmarks and leaderboards Zheng et al. (2023), the Mistral-7B-Instruct model recorded the highest hallucination density among the evaluated models.
Hallucination vs Length
We divided the data for each type of task into five equal buckets based on context/response length and calculated the average number of hallucination spans per response within each bucket. As shown in Table 4, with a few exceptions, there is a clear overall trend of an increase in the average number of hallucinations as the context length or response length grows.
Location of Hallucinations
In Figure 3, we present an analysis of the distribution of hallucinatory spans within responses. Generally, hallucinations are more likely to occur towards the end of responses, a pattern that is particularly evident in question-answering and summarization tasks. Compared to other tasks, the data-to-text task has a relatively higher occurrence of hallucinations in the first half, as indicated by the two bright areas on the heatmap.
Figure 3: Heatmaps of normalized hallucination occurrence positions. The numerical values on the colorbar represent the proportion of hallucinations occurring at a specific location relative to the total number of hallucinations in that task.
5 Experimental Setup
5.1 Hallucination Detection Algorithms
Using RAGTruth, we conducted experiments with the following four distinct algorithms for hallucination detection:
Hallucination Detection Prompt: Hallucination detection prompts are manually crafted to instruct LLMs (GPT-4-turbo and GPT-3.5-turbo) in assessing whether a given reference-response pair contains hallucinated content and to identify the corresponding hallucinated spans in the response. For detailed information about these prompts, please refer to Appendix C.
Methods Question Answering Data-to-text Summarization Overall
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
[…]
Finetuned Llama-2-13B 82.9 58.6 68.7 91.8 88.1 89.9 79.4 48.1 59.9 88.6 74.1 80.7
Table 5: The response-level hallucination detection performance for each baseline method across different tasks and different models.
Methods Question Answering Data-to-text Summarization Overall
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
[…]
Finetuned Llama-2-13B 72.3 46.6 56.6 65.3 51.4 57.6 73.5 31.1 43.7 69.1 45.4 54.8
Table 6: The span-level detection performance for each baseline method across different tasks and different models.
SelfCheckGPT (Manakul et al., 2023): SelfCheckGPT employs a zero-resource, sampling-based method to fact-check the responses of black-box models. When processing each response in RAGTruth, 5 responses from other models serve as references, and GPT-4-turbo/GPT-3.5-turbo are queried to verify consistency. By scanning each sentence in a response, the hallucination position can be determined as well.
LMvLM (Cohen et al., 2023): LMvLM is an approach that employs a multi-turn interaction between two Language Models that aim to discover inconsistencies through cross-examination.
LLM Finetuning: Llama-2-13B has been fine-tuned using the training set from RAGTruth. The model takes the context-response pair with proper instructions as the input and treats the hallucinate span as the targeted generation output. We employed LoRA Hu et al. (2021) fine-tuning with an initial learning rate of 3e-4, and limiting the training to 3 epochs, all conducted on 4 A100 GPUs.
5.2 Data Split
All detection algorithms are tested on the same RAGTruth test set, which consists of 450 instances in total, derived by randomly selecting 150 instances from each task type. The rest of the data is used to fine-tune the LLama-2-13B model, as previously mentioned.
5.3 Evaluation Metrics
Response-level Detection
We report precision, recall, and F1 score for each detection algorithm and its variants across different tasks.
Span-level Detection
We calculate the overlap between the generated span and human-labeled span and report the precision, recall, and f1 score at the char-level.
6 Experimental Results
6.1 Response-level Detection
The results in Table 5 reveal that hallucination detection remains a significant challenge in the context of RAG for all existing detection methods. Even when reference information is available, the responses generated may still include hallucinations, which current LLMs cannot reliably identify. The most advanced LLM, GPT-4-turbo, achieves only an average F1 score of 68.3%. In comparison, the widely-used GPT-3.5-turbo often fails to discern hallucinated responses, often identifying most examples as non-hallucinated, highlighting the difficulty in detecting naturally occurring hallucinations. For another notable baseline, SelfCheckGPT also shows unsatisfactory performance in this regard, achieving an average F1 score of 60.5% with GPT-4-turbo and 36.6% with GPT-3.5-turbo.
By utilizing our human-labeled high-quality training set, a fine-tuned Llama-2-13B can achieve the best performance with an average 80.7% f1 score. This shows the effectiveness of our data in improving the model’s detection ability on hallucination.
6.2 Span-level Detection
RAGTruth, as a hallucination corpus with fine-grained span labels, enables us to present experimental results for span-level detection, serving as a baseline for future research. As shown in Table 6, the overall performance of the current detection method is sub-optimal, highlighting the challenges in span-level detection. Even the advanced GPT-4-turbo tends to incorrectly classify many non-hallucinated contents with a low precision of 24%. While our fine-tuned model shows improved capability in identifying hallucinated spans by achieving an averaged f1 score of 48.4%, it still falls short of perfect detection, emphasizing the inherent difficulties of this task.
Group Treatment Valid Response Num Hallucination Rate
1 Sample From Llama-2-7B-chat 450 51.8
Mistral-7B-Instruct 450 57.6
Sampling Strategy Random selection 450 55.1(-)
Least detected spans 450 43.1($\downarrow$21.8%)
No detected spans ${}^{\dagger}$ 23.9($\downarrow$56.6%)
2 Sample From GPT-3.5-Turbo-0613 450 10.9
GPT-4-0613 450 9.3
Sampling Strategy Random selection 450 10.4(-)
Least detected spans 450 5.3($\downarrow$41.0%)
No detected spans ${}^{\dagger}$ 4.9($\downarrow$52.9%)
[…]
We also report the detection performance across four different types of hallucination spans. In the current stage, as we have not differentiated the types of detected hallucinations, we only report the char-level recall for different types of hallucinations. As indicated in Figure 4, the detection of evident hallucinations proves more effective compared to that of subtle hallucinations.
Figure 4: The span-level recalls of different models on four types of hallucinations.
6.3 Hallucination Suppression
We tested the effectiveness of hallucination suppression using our finetuned hallucination detection model. For the 450 instances in the test set, we employed two strategies to select a final output from two responses generated by two different models with similar hallucination density: one strategy is based on choosing the response with fewer predicted hallucination spans, and the other is based on the absence of detected hallucination span. When the number of hallucination spans detected in both candidate responses is the same, one will be chosen at random. Due to limited response candidates, not all instances have a response that conforms to the second strategy. In practical scenarios, this issue can be addressed by increasing the number of candidate responses. We employed random selection as a simple baseline for comparison.
The results shown in Table 7 indicate that with the help of the hallucination detector, both strategies can significantly reduce hallucination rate. For the relatively small Llama-2-7B-chat and Mistral-7B-Instruct models, compared to random selection, the first strategy reduced the hallucination rate by 21.8%, while the second strategy achieved a reduction of 56.6%. Even for models with a low hallucination rate, specifically GPT-3.5-Turbo and GPT-4, employing the finetuned hallucination detector for sampling can still further reduce the rate of hallucinations. The two strategies yielded a reduction in hallucination rates of 41% and 52.9%, respectively. These results demonstrate the potential of an efficient hallucination detection model in developing trustworthy RAG LLMs.
7 Conclusion
In this paper, we introduce RAGTruth, a large-scale corpus of naturally generated hallucinations, featuring detailed word-level annotations tailored for Retrieval-augmented Generation (RAG) scenarios. Our work includes an in-depth analysis of the interplay between hallucinations and various factors, such as task types, models being used, and contextual settings.
Additionally, we conduct empirical benchmarks of several hallucination detection approaches using our corpus. We show that fine-tuning Llama with RAGTruth leads to competitive performance in hallucination detection. This implies that by using a high-quality dataset such as RAGTruth, it is possible to develop specialized hallucination detection models that are highly effective when compared to prompt-based methods using general models such as GPT-4. The dataset paves the way for future research in this direction.
Simultaneously, our findings reveal that identifying hallucinations in RAG contexts, particularly at the span level, remains a formidable challenge, with current methods still falling short of reliable detection. We hope that RAGTruth, with its comprehensive span annotations, can assist the development of hallucination detection techniques for retrieval augmented generation.
8 Acknowledgement
We appreciate the valuable feedback and assistance from Shizhe Diao. We thank Dongfang Li for her support in illustration.
Appendix A Hallucination Examples
Evident Conflict
Original content: They’ve sent about 500 of the most seriously injured people for treatment in Kathmandu already – the majority by road.
Generative content: The team airlifted 500 injured people… to Kathmandu for medical treatment.
Subtle Conflict
Original content: This is the section that lets the government deny trademark protection to a mark that is "immoral," "scandalous" or "disparaging."
Generative content: …violates Section 2(a) of the trademark act, which prohibits marks that are "immoral, scandalous, or disqualified."
Evident Introduction of Baseless Information
Original content: The couple divorced in 1968… She had no contact with the surviving members of the band until meeting up at the 2006 Las Vegas premiere of "The Beatles Love."
Generative content: The couple divorced in 1968, but later reconciled and attended the 2006 Las Vegas premiere of "The Beatles Love."
Subtle Introduction of Baseless Information
Original content: in a restaurant review where a customer briefly mentions a "happy hour" without specifying the timing
Generative content: assumes it to be between 4-7 PM.
Table 8: Examples of the four types of hallucinations.
Appendix B Response Generation Prompts
Summarization
Summarize the following news within {word_num} words:
{news}
output:
Question answering
Answer the following question:
{question}
Bear in mind that your response should be strictly based on the following 3 passages:
{passages}
In case the passages do not contain the necessary information to answer the question, please reply with: "Unable to answer based on given passages."
Data-to-text
Instruction:
Write an objective overview about the following local business based only on the provided structured data in the JSON format.
You should include details and cover the information mentioned in the customers’ review. The overview should be 100 - 200 words. Don’t make up information.
Structured data:
{json_data}
Overview:
Table 9: Prompts for generating responses for the three types of tasks. word_num is min(200, word_num_of_news//4)
Appendix C Hallucination Detection Prompts
Summarization
Article: {context}
Summary: {sentence}
Please analyze the provided summary and indicate whether there are any hallucinated (invented or incorrect) parts. If there are, please specify the exact span that is hallucinated.
Please provide a response in the following format:
{"hallucination_detected": "yes/no",
"hallucinated_span": "If yes, the exact span of every hallucinated text part in list format; if no, leave this empty."}
Question answering
Passages: {passage}
Question: {question}
Answer: {answer}
Please analyze the provided Passages, Question, and Answer, and indicate whether there are any hallucinated (invented or incorrect) parts in the answer. If there are, please specify the exact span that is hallucinated.
Please provide a response in the following format:
{"hallucination_detected": "yes/no",
"hallucinated_span": "If yes, the exact span of every hallucinated text part in list format; if no, leave this empty."}
Data-to-text
Information: {information}
Overview: {response}
Please analyze the provided business information and its overview, and indicate whether there are any hallucinated (invented or incorrect) parts in the answer. If there are, please specify the exact span that is hallucinated.
Please provide a response in the following format:
{"hallucination_detected": "yes/no",
"hallucinated_span": "If yes, the exact span of every hallucinated text part in list format; if no, leave this empty."}
Table 10: Prompts for detecting hallucination for the three types of tasks.